# PeakWeather: MeteoSwiss Weather Station Measurements for Spatiotemporal Deep Learning

### v1

@@ 4 Machine learning tasks on PeakWeather
Wind forecasting is the application considered in the experiments below (Section 5). Accurate wind forecasts provide useful information for diverse applications such as anticipating extreme weather and optimizing renewable energy production. Forecasting wind in complex terrain, however, is challenging due to the dominant influence of local conditions. Exposure to prevailing large-scale winds and related effects, such as channeling and blocking, affects wind speeds at individual locations. In addition, distinct phenomena such as mountain valley wind systems dominate the temporal evolution.
@@
Missing data imputation addresses temporal gaps caused by missing or erroneous sensor readings, aiming to reconstruct the time series at observed locations. In contrast, virtual sensing (or spatial interpolation) focuses on estimating weather variables at locations without prior measurements using spatial patterns from nearby stations. These tasks can involve conditioning on spatially and temporally related observations (possibly from future time steps too), topographical information, and other related variables at the stations to improve accuracy. Despite the high reliability of SwissMetNet, missing observations remain inevitable (Table 2). Moreover, the continuous updates to the SwissMetNet network, where new stations are added, some are reconfigured, and others are removed, motivate considering inductive learning methods, where model predictions can extend to new, unseen stations based on their spatial and contextual characteristics.
Graph structure learning
As demonstrated in Section 5, better predictions can be issued by leveraging relational information among the time series. However, it is not always evident which station/variable would benefit from exchanging information. While it is common to extract a graph of binary relations from, e.g., physical proximity or correlations among time series, graph structure learning emerged as a task where relations can be treated as latent variables that are learned data-driven, optimizing a prediction task of interest. Given the complex Swiss terrain and the specificity of each station’s location, performing graph structure learning in PeakWeather appears sound and insightful.
@@ 5 Experiments: a use case in wind forecasting
In this section, we address a wind forecasting task, selected here as a use-case spatiotemporal modeling task to showcase the potential of PeakWeather. In particular, we show how the historical measurements from the stations and the topographical features can be integrated to yield accurate predictions, as compared to the reference NWP model predictions.
@@
We consider the task of predicting wind direction and speed for the next 24 hours. Specifically, we consider hourly data and set the wind horizontal velocity vector as target ${\bm{y}}_{t}^{i}\in\mathbb{R}^{2}$. The components of ${\bm{y}}_{t}^{i}$, also called $u$ and $v$ components, are constructed from variables wind_speed and wind_direction and indicate the eastward and northward wind components. For this task, we consider data from the 160 meteorological stations and with static attributes to encode the latitude, longitude, station height, and other topographic features. Notably, resampling from 10-minute resolution to 1-hour (while accounting for variable-specific aggregation strategies), $u$ and $v$ component extraction, and weather station filtering can all be carried out via the utilities in the PeakWeather dataset library. The considered models described in the following sections are trained on the historical data until December 31st, 2023 (i.e., seven years), tested on the last year of available data, and validated on the remaining months. A graph similar to that in Figure 1(b) is constructed by connecting stations based on their haversine distance. Further details on the training procedure, model selection, and graph construction are provided in Appendix D.
@@
We train models minimizing the Energy Score (ES) (gneiting2007strictly, )—an extension of the Continuous Ranked Probability Score (CRPS) for multivariate predictions—which evaluates the quality of probabilistic forecasts accounting for both the calibration and sharpness of predicted distributions; the ES is estimated and optimized via Monte Carlo sampling. Models are then evaluated on the test set using both probabilistic and point prediction metrics. Specifically, we report the ES for the bivariate wind velocity, and the mean absolute error on the same vector (MAE-vel) to assess point-forecast accuracy. In addition, we include the MAE for the wind speed (MAE-speed) and for the wind direction (MAE-dir), the latter computed as the mean angular distance between the target and predicted wind velocity. All metrics are reported across several lead times up to 24 hours. MAE, MAE-speed, and MAE-dir are evaluated considering the median of the wind velocity, speed and direction, respectively, over 100 Monte Carlo samples (11 samples for ICON). Missing observations are appropriately masked during evaluation.
@@ 5.1 Forecasting models and baselines
We consider two main deep learning forecasting architectures, recurrent neural networks (RNNs) and spatiotemporal graph neural networks (STGNNs) (jin2024survey, ), which we contrast with two persistence baseline models (PM) and the NWP model provided within PeakWeather.
The RNN models process each station independently from the others. It is a multilayer RNN equipped with gated recurrent units (chung2014empirical, ). The architectures feature an encoder block to aggregate the input ${\bm{x}}_{t-w}^{i}$ and exogenous variables ${\bm{u}}_{t-w}^{i}$ and ${\bm{v}}^{i}$ of station $i$ for each input time step $t-w$, before feeding them to the RNN, and a decoder responsible for making forecasts at the desired lead times; both encoder and decoder are implemented as multilayer perceptrons. The STGNN models implement a time-then-space architecture (cini2023graph, ), following the same RNN architecture above, but composing GNN layers on top of the RNN, before issuing predictions via the decoder. For both RNN and STGNN, the model outputs bidimensional vectors $\hat{\bm{y}}_{t+h}^{i}$ for each lead time $h$ as point predictions. Instead, for probabilistic forecasts, for each lead time $h$ independently, a bivariate Gaussian distribution is learned via the reparametrization trick and conditioned on the model inputs (i.e., ${\bm{x}}_{t-W:t}^{i}$, ${\bm{u}}_{t-W:t}^{i}$, and ${\bm{v}}^{i}$). Optionally, node-level learnable embeddings ${\bm{e}}^{i}\in\mathbb{R}^{D_{e}}$ are learned end-to-end alongside the rest of the model parameters and are used to condition the model encoder and decoder outputs to account for station-specific dynamics (cini2023taming, ); we discuss them in Section 5.2. Further details on the model architectures follow in the supplementary material.
Two versions of the PM are implemented: one particularly relevant for short-term, PM-st, and the other, PM-day, intended to replicate the daily patterns of the target variables. PM-st is a baseline model that predicts $\hat{\bm{y}}_{t+h}^{i}={\bm{y}}_{t-1}^{i}$ as the last available observation of the target for each lead time $h=0,\dots,H-1$, a valid assumption for very short-term forecasts. Conversely, the output of PM-day is $\hat{\bm{y}}_{t+h}^{i}={\bm{y}}_{t+h-24}^{i}$, recalling that in this specific problem instance we are considering hourly data and forecasting horizon of $H=24$. While less evident than with other weather quantities, the wind can have a diurnal seasonality, especially in mountainous valleys. For probabilistic forecasts, independent Gaussian distributions $\mathcal{N}(\hat{\bm{y}}_{t+h}^{i}[d],\sigma_{t-W:t}^{i})$ are produced for every $d,h,t$ and $i$, with $\sigma_{t-W:t}^{i}$ being the sampling standard deviation estimated on the time window ${\bm{y}}_{t-W:t}^{i}[d]$.
Finally, the ICON numerical predictions (see Section 3) provided within PeakWeather are compared with the above models; as ICON has been operational since May 2024 issuing predictions every 3 hours, an additional specific subset of the test set (NWP test set) will be considered for comparing with this model.
Sample 24-hour forecasts of trained models and ICON are shown in Appendix E.
@@ 5.2 Results
We present three sets of experiments to showcase how PeakWeather can be used and why it is a valuable resource for the deep learning and spatiotemporal modeling research communities. (i) We analyze the impact of incorporating station-specific information—a key feature of PeakWeather—highlighting the presence of local effects. (ii) We assess the wind prediction accuracy of different models, showcasing the competitiveness of graph deep learning models, like STGNNs, for spatiotemporal modeling, and providing baseline prediction accuracies for future research. Finally, (iii) we compare the best-performing models with forecasts from the ICON numerical weather prediction (NWP) model, demonstrating the challenging nature of the task.
Table 3: Assessment of the impact of station-specific features on the prediction performance. Results are aggregated over five runs with different random seeds and are reported as mean $\pm$ standard deviation.
Model MAE-vel ES
RNN 1.173$\pm$0.005 1.334$\pm$0.008
RNN(${\bm{v}}$) 1.134$\pm$0.006 1.281$\pm$0.007
RNN(${\bm{e}}$) 1.143$\pm$0.006 1.292$\pm$0.009
RNN(${\bm{e}},{\bm{v}}$) 1.133$\pm$0.002 1.279$\pm$0.002
STGNN 1.096$\pm$0.001 1.241$\pm$0.004
STGNN(${\bm{v}}$) 1.088$\pm$0.004 1.227$\pm$0.004
STGNN(${\bm{e}}$) 1.094$\pm$0.003 1.233$\pm$0.001
STGNN(${\bm{e}},{\bm{v}}$) 1.094$\pm$0.006 1.236$\pm$0.011
@@
Table 3 presents the test performance for RNN and STGNN models trained with and without station-specific attributes: static vectors ${\bm{v}}^{i}$ of station descriptors and learnable node embeddings ${\bm{e}}^{i}$. For RNNs, incorporating either ${\bm{v}}^{i}$, ${\bm{e}}^{i}$, or both leads to substantial improvement in the predictive accuracy, highlighting the importance of station-level information in the absence of built-in spatial modeling. For STGNNs, however, the inclusion of station-specific features results in only marginal performance differences. This suggests that STGNNs already capture key spatial dependencies through message passing, making explicit station embeddings less critical. Nevertheless, STGNNs consistently outperform RNNs overall, confirming their advantage in spatiotemporal modeling.
Figure 2: Performance comparison of deep learning models with ICON and persistence model baselines for wind forecasting. Results are averaged over five runs with different random seeds, except for ICON. As the persistence models require no training, their variability arises solely from Monte Carlo sampling during metric evaluation. Shaded areas indicate $\pm$3 standard deviations.
@@
The top part of Figure 2 shows forecasting performance across lead times for all models. All trainable models include static station attributes; among them, RNN and STGNN also utilize learnable node embeddings, while RNN-noemb is identical to RNN except for the omitted node embedding ${\bm{e}}^{i}$ and discarded static exogenous ${\bm{v}}^{i}$. PM-st baseline provides a strong reference, particularly in the short term. As the horizon increases, PM-day marginally outperforms PM-st, although the difference is minor within the considered 24 hours; as expected, PM-st and PM-day coincide at 24h predictions. All deep learning models substantially outperform the PM baselines across all horizons, both in terms of MAE and probabilistic accuracy, measured by the ES metric. Among them, the STGNN model achieves the best overall performance, underscoring the value of incorporating spatial structure. As expected, performance degrades with increasing forecast horizon. This trend holds for all models, with the exception of the persistence baselines: PM-day effectively remains fixed at 24h forecasting horizon, while PM-st benefits from daily wind patterns. Notably, wind direction is consistently harder to predict than wind speed as evidenced by smaller improvements over the PM baselines.
Comparison with ICON NWP forecasts
Figure 2 also includes model performance on the NWP test set, the subset of the full 1-year test set aligned with ICON forecasts. First, we note that the forecasting metrics on this subset are consistent with those observed on the full test set. Then, we observe that ICON demonstrates relatively stable performance across all lead times, including short-term forecasts. However, its accuracy does not improve significantly at near-term horizons; this is likely due to spatial mismatches between the gridded NWP output and the fine-scale conditions at station locations and consequent challenges in assimilating such local information in the model. Comparing ICON to the deep learning models reveals that STGNN delivers more accurate predictions in the nowcasting regime, and up to around 12 hours ahead. For longer horizons, and particularly for wind direction, ICON outperforms the data-driven approaches, suggesting that physics-based models still hold an advantage in capturing larger-scale dynamics over extended periods.
@@ 6 Conclusion
The paper introduces PeakWeather, a high-resolution dataset of validated ground-based weather measurements collected from the SwissMetNet network operated by MeteoSwiss. Designed to support and accelerate research in spatiotemporal deep learning, the dataset provides dense and diverse weather observations from 302 stations distributed across Switzerland. The dataset includes eight meteorological variables sampled every 10 minutes for more than eight years. Notably, the data is enriched with topographical descriptors to account for the complex Swiss terrain and with ensemble forecasts from a state-of-the-art NWP model currently operational at MeteoSwiss for benchmarking.
Beyond its scale and quality, PeakWeather distinguishes itself by supporting a wide range of machine learning tasks. These include time series forecasting, virtual sensing (spatial interpolation), graph structure learning, and missing data imputation. PeakWeather data is provided alongside a framework-agnostic library that offers data in ready-to-use tensor formats for seamless integration into modern deep learning workflows. Experiments on wind forecasting conclude the paper, showcasing the timeliness of PeakWeather, its relevant features for spatiotemporal modeling and reinforcing the emerging role of deep learning as a complementary tool to traditional NWP systems. We are confident that PeakWeather will foster both fundamental machine learning research as well as advancing data-driven approaches in meteorological modeling.

### v2

@@ 4 Machine learning tasks on PeakWeather
Wind forecasting is the main application considered in the experiments in Section 5. Accurate wind forecasts provide useful information for diverse applications such as severe weather and renewable energy management. Forecasting wind in complex terrain, however, is challenging due to the dominant influence of local conditions. Exposure to prevailing large-scale winds and related effects such as channeling and blocking affect wind speeds at individual locations. In addition, distinct phenomena such as mountain valley wind systems dominate the temporal evolution. Temperature forecasting is studied in Appendix F.
Local effects and node-specific components
As anticipated in (1), time series observed at different locations may exhibit local effects, namely variations that are specific to individual locations or small regions. Such specificity can arise from, e.g., unobserved factors and station-specific characteristics. In the case of PeakWeather, the complex Swiss terrain in which measurement stations are deployed plays a role. Accounting for local effects poses a significant modeling challenge, as it requires balancing the extraction of global patterns common to all time series with the ability to specialize predictions to individual ones, under a less favorable sample size [25]. In the experimental section below, we demonstrate the importance of incorporating station-specific components for achieving accurate predictions. In particular, we show the positive impact of conditioning predictions on station-level topographic descriptors ${\bm{v}}^{i}$ as well as on learnable station-specific embeddings ${\bm{e}}^{i}\in\mathbb{R}^{D_{e}}$.
Relational inductive biases for time series processing
Beyond local components, time series processing can be significantly improved by exploiting the underlying (functional) dependencies among series acquired from spatially distributed stations. Rather than modeling each sequence in isolation, one can incorporate relational information as an inductive bias, allowing models to condition predictions on related time series and reduce the risk of overfitting to local, spurious patterns. These dependencies can be encoded via a graph structure represented by an adjacency matrix ${\bm{A}}\in\{0,1\}^{N\times N}$, where each node corresponds to a station and non-zero entries denote pairwise relationships. When needed, $\bm{A}$} may be enriched with edge weights or attributes reflecting the strength or type of interaction (e.g., distance, elevation difference, empirical correlation). This setup is at the core of graph deep learning approaches for time series processing [26], where message-passing operators like graph neural networks (GNN) are exploited to exchange information across the graph topology. In the context of meteorological data, such relations are often assumed to be time-invariant. However, changes to the sensor network—such as the addition, removal, or relocation of stations—may require considering a dynamic or evolving graph. While spatial proximity naturally induces correlation (i.e., signal smoothness), this assumption breaks down at larger scales or in complex terrains. In alpine regions, for instance, valleys, ridges, and altitude gradients introduce significant variability, making long-range dependencies non-trivial to infer. Particularly in datasets like PeakWeather, this motivates considering relations learned in a data-driven manner and tailored to the given task, as discussed in the next paragraph.
Graph Structure Learning
Better predictions can be obtained by leveraging relational information among the time series, as demonstrated in the next Section 5. However, it is not always evident which stations/variables would benefit from exchanging information. While it is common to extract a graph of binary relations from, e.g., physical proximity, correlations among time series, graph structure learning has emerged as a task where relations can be treated as latent variables that are learned in a data-driven manner, often optimizing a prediction task of interest. Given the complex Swiss terrain and the specificity of each station’s location, performing graph structure learning in PeakWeather appears sound and insightful.
@@
Another challenge posed by real-world data concerns temporal gaps in the measurements and unavailable observations at certain locations. A related task is broadly known as spatial interpolation. It aims to estimate unmeasured weather variables by leveraging spatial patterns learned from nearby stations, topographical features, and, when available, related variables at the same location. While missing data imputation focuses on filling temporal gaps (missing or erroneous data), virtual sensing refers to predicting variables at locations without any prior measurements. Notably, addressing these tasks may involve conditioning predictions on observations at future time steps. Despite the high reliability of SwissMetNet, missing observations remain inevitable (Table 2). Moreover, the continuous updates to the SwissMetNet network, where new stations are added, some are reconfigured, and others are removed, motivate considering inductive learning methods, where model predictions can extend to new, unseen stations based on their spatial and contextual characteristics. In Appendix F, we experiment on this inductive setting considering the temperature as target variable.
@@ 5 Experiments: a use case in wind forecasting
In this section, we consider wind forecasting as a representative spatiotemporal modeling task to showcase the potential and breadth of PeakWeather. We evaluate the capabilities and advantages of deep spatiotemporal models in comparison to the reference NWP model based on atmospheric simulations. Alongside forecasting, we demonstrate the versatility of PeakWeather to study different related learning challenges, including the modeling of station-specific effects and the learning of relational information directly from data. In particular, we show the positive impact of incorporating topographical descriptors, learnable station-specific embeddings, and graph structure learning.
@@
We consider the task of predicting wind direction and speed for the next 24 hours. Specifically, we consider hourly data and set the wind horizontal velocity vector as target ${\bm{y}}_{t}^{i}\in\mathbb{R}^{2}$. The components of ${\bm{y}}_{t}^{i}$, also called $u$ and $v$ components, are constructed from variables wind_speed and wind_direction and indicate the eastward and northward wind components. For this task, we consider data from the 160 meteorological stations and with static attributes encoding the latitude, longitude, station height, and other topographic features. Notably, resampling from 10-minute resolution to 1-hour (while accounting for variable-specific aggregation strategies), $u$ and $v$ component extraction, and weather station filtering can all be carried out via the utilities in the PeakWeather dataset library. The considered models described in the following sections are trained on the historical data before October 2024 (i.e., almost eight years), tested on the last year of available data, and validated on a held-out set of training data from October 2023 to September 2024. A graph similar to that in Figure 3(b) is constructed by connecting stations based on their horizontal Euclidean distance; the graph is passed as input to all models requiring one. Further details on the training procedure, model selection, and graph construction are provided in Appendix D.
@@
We train models minimizing the Energy Score (ES) [27]—an extension of the Continuous Ranked Probability Score (CRPS) for multivariate predictions—which evaluates the quality of probabilistic forecasts accounting for both the calibration and sharpness of predicted distributions; the ES is estimated and optimized via Monte Carlo sampling. Models are then evaluated on the test set using both probabilistic and point prediction metrics. Specifically, we report the ES for the bivariate wind velocity, and the mean absolute error on the same vector (MAE Velocity) to assess point-forecast accuracy. In addition, we include the MAE for the wind speed (MAE Speed) and for the wind direction (MAE Direction), the latter computed as the mean angular distance between the target and predicted wind velocity. All metrics are reported across several lead times up to 24 hours. The metrics are evaluated considering 100 Monte Carlo samples (11 samples for ICON), with the MAEs based on the sample median. Missing observations are appropriately masked during evaluation.
@@ 5.1 Forecasting models and baselines
We consider two main families of deep learning forecasting architectures, temporal models like recurrent neural networks and spatiotemporal models based on message passing and graph neural networks [28].
The temporal models RNN and TCN are multilayer networks processing each station independently from the others. The architectures feature an encoder block to aggregate the input ${\bm{x}}_{t-w}^{i}$ and exogenous variables ${\bm{u}}_{t-w}^{i}$ and ${\bm{v}}^{i}$, before feeding them to the RNN/TCN temporal processing, and a decoder responsible for making forecasts at the desired lead times; both encoder and decoder are implemented as multilayer perceptrons. The spatiotemporal models are MP-RNN and MP-TCN and implement a time-then-space architecture [26], following the same high-level architecture of the temporal models, but composing GNN layers on top of the temporal processing, before issuing predictions via the decoder.
As competitive models from the literature, we consider spatiotemporal models DCRNN [29] and AGCRN [30], based on RNNs, and GraphWaveNet [31], based on temporal convolution. As temporal models, we include the foundation model Chronos-2 [32], to reflect the growing relevance of large pretrained models in time series forecasting.
For reference, two versions of the persistence model (PM) are implemented: one particularly relevant for short-term, PM-st, and the other, PM-day, intended to replicate the daily patterns of the target variables. PM-st is a baseline model that predicts $\hat{\bm{y}}_{t+h}^{i}={\bm{y}}_{t-1}^{i}$ as the last available observation of the target for each lead time $h=0,\dots,H-1$, a valid assumption for very short-term forecasts. Conversely, the output of PM-day is $\hat{\bm{y}}_{t+h}^{i}={\bm{y}}_{t+h-24}^{i}$, recalling that in this specific problem instance we are considering hourly data and a forecasting horizon of $H=24$. While less evident than with other weather quantities, the wind can have a diurnal seasonality, especially in mountainous valleys.
Finally, the ICON numerical predictions (Section 3) provided within PeakWeather and obtained from ICON-CH1-EPS are compared with the above models.
@@ 5.2 Results
We present four sets of experiments to illustrate how PeakWeather can be used in practice and to highlight its value for the deep learning and spatiotemporal modeling research communities. (i) We investigate the impact of incorporating station-specific information—a core feature of PeakWeather—highlighting the presence of local effects in the data. (ii) We evaluate the benefits of graph-based models and of learning relational information directly from the data. (iii) We benchmark the wind prediction accuracy of a range of models, showcasing the competitiveness of graph deep learning approaches for spatiotemporal modeling, and providing reference prediction accuracies for future research. Finally, (iv) we compare the best-performing models against forecasts from the NWP model ICON, demonstrating the challenging nature of the task and highlighting the potential held by deep learning methods in this field.
Table 3: Assessment of the impact on the prediction performance of including station-specific features and learning a graph structure from data. The family of models based on temporal convolution is considered here; for other models see Table 4. The first columns indicate whether the corresponding model relies on station-specific topographic descriptors ${\bm{v}}^{i}$ (Topo.), learnable node embeddings ${\bm{e}}^{i}$ (Emb.), an input graph (In.G.), and a graph structure learned from the data (GSL). Results are aggregated over five runs with different random seeds and are reported as mean $\pm$ standard deviation. For both metrics, the lower the better. Symbol “-” indicates non-applicability.
TCN Family
Model Components Metrics
Topo. Emb. In.G. GSL MAE Vel. En. Score
[…]
@@
Table 3 reports test performance for temporal and spatiotemporal models trained with and without station-specific attributes, namely static station descriptors ${\bm{v}}^{i}$ and node-level learnable embeddings ${\bm{e}}^{i}\in\mathbb{R}^{D_{e}}$. Unlike the topographic features in ${\bm{v}}^{i}$, the embeddings ${\bm{e}}^{i}$ are learned end-to-end together with the other model parameters and then used to condition both the encoder and decoder. For purely temporal models, incorporating either ${\bm{v}}^{i}$, ${\bm{e}}^{i}$, or both leads to substantial improvements in accuracy, highlighting the importance of station-level information. Spatiotemporal models also benefit from the inclusion of station-specific features, although the gain is less pronounced. This suggests that message passing already provides a degree of localization by conditioning predictions on related time series, thereby reducing the contribution of station embeddings. Nevertheless, spatiotemporal models consistently outperform temporal ones overall.
Relevance of graph structure learning
Table 3 further shows that replacing or augmenting the (heuristic) proximity-based graph with a data-driven graph yields significant improvements in prediction accuracy. This learned relational structure is jointly optimized with the remaining model parameters on the downstream task.
Similar trends are observed across RNN, MP-RNN, and GraphWaveNet architectures that we report in Table 4. Overall, these results emphasize the importance of accounting for both station-specific features and relational information, and demonstrate that such structure can be effectively learned from data, tailored to the considered task.
Figure 2: Performance comparison of deep learning models with ICON and persistence model baselines for wind forecasting. Results are averaged over five runs with different random seeds, except for ICON and Chronos-2. Shaded areas indicate $\pm$3 standard deviations. The persistence models, ICON and Chronos-2 require no training. The variability of the persistence models and ICON arises solely from Monte Carlo sampling during evaluation and the ensemble forecasts, respectively. Chronos-2 provides quantile predictions of the velocity, which are used to compute the speed and direction.
@@
Figure 2 shows forecasting performance across lead times for selected models. All trained models account for topographic attributes ${\bm{v}}^{i}$ and learnable embeddings ${\bm{e}}^{i}$. While MP-TCN operates on the proximity graph, MP-TCN-gsl replaces it with a learned one. PM-st prediction accuracy provides reference performance, particularly relevant in the short term. As the horizon increases, PM-day marginally outperforms PM-st, although the difference is minor within the considered 24 hours; as expected, PM-st and PM-day coincide at 24h predictions. All deep learning models substantially outperform the PM baselines across all horizons, both in terms of MAE and probabilistic accuracy, measured by the ES metric. Among them, the spatiotemporal models achieve the best overall performance, underscoring the value of incorporating spatial structure; these results are also confirmed by Table 5 in Appendix E. As expected, performance degrades with increasing forecast horizon. This trend holds for all models, with the exception of the persistence baselines: PM-day effectively remains fixed at 24h forecasting horizon, while PM-st benefits from daily wind patterns. Notably, wind direction is consistently harder to predict than wind speed as evidenced by the smaller improvements over the PM baselines. Similar results abserved for temperature forecasting in Appendix F.
Comparison with ICON
We observe that ICON demonstrates relatively stable performance across all lead times, including short-term forecasts. However, its accuracy does not improve significantly at near-term horizons; this is likely due to spatial mismatches between the gridded NWP output and the fine-scale conditions at station locations and consequent challenges in assimilating such local information in the model. Comparing ICON to the deep learning models reveals that the MP-TCN deliver more accurate predictions in the nowcasting regime, and up to approximately 12 hours ahead. For longer horizons, and particularly for wind direction, ICON outperforms the data-driven approaches, suggesting that physics-based models still hold an advantage in capturing larger-scale dynamics over extended periods. Notably, the DL models are more competitive when assessed with the Energy Score, as also visible in Figure 8. This is likely due to the fact that data-driven models better represent the local specifics at station measurement locations.
@@ 6 Conclusion
The paper introduces PeakWeather, a high-resolution dataset of validated ground-based weather measurements collected from the SwissMetNet network operated by MeteoSwiss. Designed to support and accelerate research in spatiotemporal deep learning, the dataset provides dense and diverse weather observations from 302 stations distributed across Switzerland. The dataset includes eight meteorological variables sampled every 10 minutes for more than eight years. Notably, the data are enriched with topographical descriptors to account for the complex Swiss terrain and with ensemble forecasts from a state-of-the-art NWP model that is currently operational at MeteoSwiss for benchmarking.
Beyond its scale and quality, PeakWeather distinguishes itself by supporting a wide range of machine learning tasks. These include time series forecasting, virtual sensing (spatial interpolation), graph structure learning, and missing data imputation. PeakWeather data are provided alongside a framework-agnostic library that offers data in ready-to-use tensor formats for seamless integration into modern deep learning workflows. Experiments on wind forecasting conclude the paper, showcasing the timeliness of PeakWeather, its relevant features for spatiotemporal modeling, and reinforcing the emerging role of deep learning as a complementary tool to traditional NWP systems. We are confident that PeakWeather will foster both fundamental machine learning research as well as advance data-driven approaches in meteorological modeling.